5r+3rh=16h. what is the equation if you solve for h?
Answer by rfadrogane(214)

5r+3rh=16h. what is the equation if you solve for h?
Sol'n:
16h-3rh = 5r, factor out the h
h(16-3r)= 5r
h = 5r/(16-3r)--answer